solve through (2,-5), parallel to 2x - y = -4
Answer by Cromlix(4381)

adjust equation into form:
y = mx + c
2x - y = -4
-y = -2x - 4
Multiply through by -1
y = 2x + 4

The gradient (m) = 2
Two lines parallel to one another
have the same gradient (m)
m1 = m2
Using the formula:
y - b = m(x - a) (m = 2) Coords (2, -5)
y - (-5) = 2(x - 2)
y + 5 = 2(x - 2)
y = 2x - 4 - 5
y = 2x - 9
Hope this helps.
:-)